Lewis and Clark Camping Supplies Inc. is borrowing $42,000 from Western State Bank. The total interest is $17,500. The loan will be paid by making equal monthly payments for the next three years.  
  
What is the effective rate of interest on this installment loan? (Input your answer as a percent rounded to 2 decimal places.)
Answer #1

3 B 1 Principal 2 Total interest 3 Nper 4 payment 5 PV 6 FV 7 Rate 8 Effective rate of interest $96,000.00 $ 14,200.00 36 306B 2 3 B C с 1 Principal $96,000.00 2 Total interest $14,200.00 Nper =3*12 4 payment =(C1+C2)C3 5 PV =C1 6 FV 0 Rate =RATE(C3,

answered by: ANURANJAN SARSAM